1) Find the Q-point value
2) Find the minimum power rating of transistor. Given : ß = 50, Vcc = -12V, Rc = 1.8kΩ, R1= 33kΩ, R2 = 5.6kΩ, RE = 560Ω.
1
Expert's answer
2020-10-05T16:03:04-0400

i)To find Q-value we need to calculate,

"V_{BB}=V_{cc}\frac{R_2}{R_1+R_2}"


=-12"\times\frac{5.6}{5.6+33}"

=-1.74V

Resistance

"R_B=R_1||R_2"


="\frac{R_1R_2}{R_1+R_2}"

="\frac{5.6\times 33}{5.6+33}"

=4.78K"\Omega"


"I_{E}=\frac{V_{BB}-V_{CC}}{\frac{R_B}{\beta+1}+R_E}"


"=\frac{-1.74-(-12)}{\frac{4.78}{51}+0.56}"


="\frac{10.26}{0.653}"

=15.71mA


"V_{CE}" ="V_{CC}-I_E(R_C+R_E)"


="-12-15.71(1.8+0.56)"

=-12-37.0756

=-49.07V

Hence Q point is (-49.07V,15.71mA)

ii) Power of a transistor is given by

P"=V_{CE}I_{CE}+V_{BE}I_{B}"


=V"_{CE}I_{CE}"

="-49.07\times 15.71"

=-0.77 Watt